If we test the following: H0: μ = 17
vs. H1: μ ≠ 17 and the test statistic (tobs.) is -2.93 for n = 16, so the p-value for this test is


Select one:
to. .01 <value p <.02
b. .02 <value p <.05
c. .02 <value p <.01
d. 0.0034
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is ,

H0 :   = 17

Ha :    17

t = -2.93

Test statistic = -2.93

n = 16

degrees of freedom = n - 1 = 16 - 1 = 15

P-value = 0.0103